solve by the substitution method
5x+2y=-19
-5x+y=43

Answer by robertb(5830)

From bottom equation, y - 43 = 5x
==> (y - 43) + 2y = -19
==> 3y - 43 = -19
==> 3y = 24
==> y = 8
==> y - 43 = 8 - 43 = -35 = 5x
==> x = -7.